### v1

# Cohomology ring of non-compact abelian arrangements

Abstract.
We give a Orlik-Solomon type presentation for the cohomology ring of arrangements in a non-compact abelian Lie group. The new insight consists in comparing arrangements in different abelian groups. Our work is based on the Varchenko-Gelfand ring for real hyperplane arrangements and from that we deduce the cohomology rings of all other abelian arrangements. As by-product, we obtain a new proof of the Orlik-Solomon relations and De Concini-Procesi ones.
1. Introduction
The topology of hyperplane arrangements is a classical subject whose study goes back to the sixties. It is originally motivated by the relations with braid groups [Del72] and configuration spaces [CLM76]. The study of arrangements was also motivated by combinatorics, in particular matroid theory, and by singularity theory. In the last few years, the new development in matroid theory renewed the interest in the topology of arrangements and their compactifications: June Huh proved inequalities for invariants of realizable matroids by using properties of hyperplane arrangements [Huh12], that were later generalized to all matroids [AHK18]. Recently, hyperplane arrangements were used as local model in algebraic geometry, especially in moduli problems [dCHM21, MMP23].
In the 2000s, De Concini and Procesi [DCP05] started the study of more general arrangements – called toric arrangements – motivated by the application to knapsack problem. Finally, abelian arrangements were introduced in 2016 by Bibby [Bib16].
An important focus on this subject is the description of the cohomology ring of the arrangement complement. Arnold identified relations between differential forms in the cohomology ring of braid hyperplane arrangements [Arn69] and Brieskorn described the cohomology ring of braid arrangements [Bri72]. A milestone is the famous result by Orlik and Solomon [OS80]: they presented the cohomology ring of complex hyperplane arrangements by generators and relations. Varchenko and Gelfand studied the cohomology ring of real hyperplane arrangements using Heaviside functions as generators [VG87]. They observed an analogy between the real and the complex case:
The relation (8)  [^1] Equation (8) of this article corresponds to relation (2) in [VG87] is the even analogue of a relation of Orlik and Solomon [OS80] for differential forms.
However, they do not find a geometric connection between the two cases.
The cohomology of subspace arrangements was determined in [GM12] as a module and a rational model was provided in [DCP95]. The multiplicative structure of the integral cohomology ring was later provided in [FZ00] and in [dLS01].
Proudfoot [Pro06] relates the VG-algebra and the OS-algebra with $\mathbb{F}_{2}$-coefficients overcoming the non-commutativity of the OS algebra. Moseley [Mos17] shows an isomorphism between the VG-algebra and the cohomology of (subspace) arrangements in $(\mathbb{R}^{3})^{n}$. The Varchenko-Gelfand presentation was generalized to oriented matroids and conditional oriented matroids [GR89, Cor02, DB23, DBPW22].
A Orlik-Solomon type presentation for toric arrangements was obtained by De Concini and Procesi [DCP05] in the unimodular case and later extended to all toric arrangements [CD17, ${}^{+}$20]. The problem of understanding arrangements of subtori was approached in a sequence of articles [DCG18, DCG19, DCGP20, DCG21, MP22].
In the more general case of abelian arrangements only few results are known: Bibby computed the Euler characteristic in the case of complex groups [Bib16] and Liu, Tran, and Yoshinaga described the additive structure of the cohomology in the case of non-compact abelian groups [LTY21].
Our main result is a new and uniform approach to the cohomology ring of non-compact abelian arrangements. Our technique consists in the pushforward of the cohomological relations from the real hyperplane case (see [VG87]) to the general non-compact abelian arrangement. The main theorem (Theorem 5.11) is an Orlik-Solomon type presentation of the cohomology ring of non-compact abelian arrangements.
We obtain a new proof of the Orlik-Solomon result [OS80] for complexified hyperplane arrangements and of the De Concini-Procesi presentation [DCP05] for unimodular toric arrangements. In all other cases our result is original.
We start by setting the notations in Section 2. In Section 3, we explicitly describe the pushforward along the map $i\colon M^{\mathbb{R}}(\mathcal{A})\to M^{a,b}(\mathcal{A})$ induced by the inclusion $\mathbb{R}\hookrightarrow\mathbb{R}^{b}\times(S^{1})^{a}$. We manipulate the Varchenko-Gelfand relations to obtain relations in the cohomology ring of central and unimodular arrangements (see Lemma 3.12). Section 4 extends the results of [LTY21] to the case of non-central arrangements (see Theorem 4.1). Then, in Section 5 we use the technique of separating coverings developed in [${}^{+}$20] to obtain the main theorem. Finally, in Section 6 we apply our theorem to the arrangements of type $A_{n}$ improving a result of [CT78] on configuration spaces in $\mathbb{R}^{b}\times(S^{1})^{a}$.
2. Definitions and notations
Consider the abelian group $G=\mathbb{R}^{b}\times(S^{1})^{a}$ canonically oriented and let $e=(0,\dots,0,1,\dots,1)$ be the unit of $G$. We denote $d:=\operatorname{dim}(G)-1=a+b-1$.
Let $\Lambda\cong\mathbb{Z}^{r}$ be a lattice with the choice of an orientation, so that $G^{r}$ has a natural orientation. A primitive element $\chi\in\Lambda$ corresponds to a group morphism $\chi\colon G^{\Lambda}\rightarrow G$. Let $E$ be a finite set with a total ordering. An abelian arrangement is a finite collection of connected subvarieties in $G^{\Lambda}$ defined by a finite list of functions $\chi_{i}$ and elements $g_{i}\in G$ indexed by $E$:
$\mathcal{A}=\{H_{i}:=\chi_{i}^{-1}(g_{i})\}_{i\in E}.$
This definition, specialized to the case $G=\mathbb{C}$, corresponds to the definition of complexified hyperplane arrangements in the central case; in the non central case we allow also complex translations.
The poset of layers (or intersection poset) of $\mathcal{A}$, $\mathcal{L}(\mathcal{A})$, is defined as the set of all connected components of intersections of elements in $\mathcal{A}$, ordered by reverse inclusion. An arrangement is called central if the intersection of all the subvarieties is non-empty, and it is called unimodular whenever all the possible intersections of elements in $\mathcal{A}$ is either connected or empty.
The complement of the arrangement is:
$M(\mathcal{A})=G^{\Lambda}\setminus\bigcup_{i\in E}H_{i}.$
We also denote $M(\mathcal{A})$ by $M^{a,b}(\mathcal{A})$, when we need to specify what abelian group $G=\mathbb{R}^{b}\times(S^{1})^{a}$ we are working with. When $G=\mathbb{R}$, we write $M^{\mathbb{R}}(\mathcal{A})$ instead of $M^{0,1}(\mathcal{A})$.
In order to define some classes in $H^{a+b-1}(G\backslash\{e\})$ we consider the immersion of a small sphere $i\colon S^{a+b-1}\hookrightarrow G\setminus e$ centered in $e$. Let $W$ be the top dimensional class generating $H_{a+b-1}(S^{a+b-1})$. We denote by $\omega$ the class in $H^{a+b-1}(G\setminus\{e\})$ dual to $i_{*}(W)$ with the orientation given by the outer normal first rule.
For $j=1,\cdots,a$, let $Y^{j}$ be the standard generators of $H_{1}(S^{1})$ and let $Y$ be the top class in $H_{a}((S^{1})^{a})$. For $x\in\mathbb{R}^{b}\setminus\{0\}$ we consider the immersion
$i_{x}\colon\quad(S^{1})^{a}$ $\hookrightarrow G\setminus\{e\}=(\mathbb{R}^{b}\times(S^{1})^{a})\setminus(0,1,\cdots,1)$
$z$ $\mapsto(x,z),$
Up to homotopy, $i_{x}$ only depends on the connected component of $\mathbb{R}^{b}\setminus\{0\}$ $x$ belongs to. In the case $b=1$ we choose $x<0$, otherwise the choice of the point does not matter. Denote by $\psi^{j}$ the class in $H^{1}(G\setminus\{e\})$ dual to $(i_{x})_{*}(Y^{j})$ and by $\psi$ the class in $H^{a}(G\setminus\{e\})$ dual to $(i_{x})_{*}(Y)$.
For any $i\in E$, the map $\chi_{i}$ restricts to $(\chi_{i}-g_{i})|_{M^{a,b}(\mathcal{A})}\colon M^{a,b}(\mathcal{A})\rightarrow G\setminus\{e\}$. So we set
$\omega_{i}=(\chi_{i}-g_{i}|_{M^{a,b}(\mathcal{A})})^{*}(\omega)\in H^{a+b-1}(M^{a,b}(\mathcal{A})),$
$\psi_{i}^{j}=(\chi_{i}|_{M^{a,b}_{\mathcal{A}}})^{*}(\psi^{j})\in H^{1}(M^{a,b}(\mathcal{A})),$
$\psi_{i}=(\chi_{i}|_{M^{a,b}_{\mathcal{A}}})^{*}(\psi)\in H^{a}(M^{a,b}(\mathcal{A})).$
Note that, when $b=1$, the classes $\psi_{i}$’s have degree $a=a+b-1$, the same degree of the elements $\omega_{i}$’s.
3. Central and unimodular arrangements
Consider a proper map $f\colon N\to M$ between oriented manifolds, it induces a pushforward map in cohomology $f_{*}$ obtained from the pushforward in Borel-Moore homology (for a general reference [HR96]) by composing with the Poincaré duality isomorphism.
{tikzcd}
The map $f_{*}$ increases the degree by $\dim M-\dim N$ and has the following properties:
  1. (1)
    (Functoriality) $f_{*}\circ g_{*}=(f\circ g)_{*}$.
  2. (2)
    (Projection formula) $f_{*}(f^{*}(y)\smile x)=y\smile f_{*}(x)$ for any $x\in H^{*}(N)$ and any $y\in H^{*}(M)$.
  3. (3)
    (Naturality) For any pullback diagram
    {tikzcd}
    with $f$ and $f^{\prime}$ proper maps and $\dim M-\dim N=\dim M^{\prime}-\dim N^{\prime}$, we have
    $g^{*}\circ f_{*}=f^{\prime}_{*}\circ h^{*}.$
  4. (4)
    (Embedding) If $f$ is a closed embedding then $f_{*}$ is the composition of the Thom isomorphism for the normal bundle
    $f_{*}\colon H^{*}(N)\xrightarrow{\operatorname{Th}}H^{*+d}(T,T\setminus N)\to H^{*+d}(M,M\setminus N)\to H^{*+d}(M)$
    where we identify the normal bundle with a tubular neighborhood $N\subseteq T\subseteq M$ and $d=\dim M-\dim N$. The second map is provided by excision theorem of $M\setminus T$ and the last one is the map from the long exact sequence of the pair $(M,M\setminus N)$.
We consider the pushforward in cohomology $i_{*}\colon H^{0}(\mathbb{R}\setminus\{0\})\rightarrow H^{a+b-1}(G\setminus\{e\})$ induced by the closed inclusion $i\colon\mathbb{R}\setminus\{0\}\rightarrow G\setminus\{e\}$, $i(x)=(x,0,\cdots,0,1,\cdots,1)$. Let $w^{+}$ and $w^{-}$ be the two standard generators of $H^{0}(\mathbb{R}\setminus\{0\})$. Poincaré Duality maps $w^{+}$ and $w^{-}$ to the classes in $H^{\operatorname{BM}}_{1}(\mathbb{R}\setminus\{0\})$ represented by the infinite chains $c^{+}$ and $c^{-}$, corresponding respectively to the positive and negative semi-axes. By checking intersection numbers, we obtain $\operatorname{PD}(\omega)=j_{*}c^{+}$ and, when $b=1$, $\operatorname{PD}(\psi)=j_{*}c^{+}+j_{*}c^{-}=j_{*}1$. Note that, if $b>1$, $j_{*}c^{+}+j_{*}c^{-}=0$ in $H^{\operatorname{BM}}_{1}(G\setminus\{e\})$. Summarizing:
$i_{*}(w^{+})=\omega,\quad i_{*}(w^{-})=\begin{cases}\psi-\omega&\text{if }b=1,\\ -\omega&\text{otherwise}\end{cases},\quad i_{*}(1)=\begin{cases}\psi&\text{if }b=1,\\ 0&\text{otherwise.}\end{cases}$ (1)
{tikzpicture}
Figure 1. In the case $(a,b)=(1,1)$, the homology classes $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}Y$ in red and $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}W$ in blue. The pushforward of the Borel Moore homology classes $c^{+},c^{-}\in H_{1}^{\operatorname{BM}}(\mathbb{R}\setminus\{0\})$.
Remark 3.1.
Consider a central arrangement $\mathcal{A}$, the linear relations among the elements $\chi_{i}\in\Lambda\simeq\mathbb{Z}^{r}$ define a representable matroid. For any subvariety $H_{i}=\ker(\chi_{i})=\ker(-\chi_{i})$ we choose one of the two descriptions. These choices determine an oriented matroid.
We will denote the circuits of the matroid by $C$ and the oriented circuits by $C=C^{+}\sqcup C^{-}$.
Now we consider $\mathcal{A}$ as a real arrangement, for any $i\in E$ let $w_{i}^{+}=\chi_{i}^{*}(w^{+})$ and $w_{i}^{-}=\chi_{i}^{*}(w^{-})$ be the classes in $H^{0}(M^{\mathbb{R}}(\mathcal{A}))$ corresponding to the inverse image of $\mathbb{R}_{+}$ and $\mathbb{R}_{-}$ through $\chi_{i}\colon\mathbb{R}^{r}\rightarrow\mathbb{R}$.
Theorem 3.2 ([VG87, Theorem 5]).
Let $\mathcal{A}$ be a central real arrangement, the ring $H^{0}(M^{\mathbb{R}}(\mathcal{A}))$ is generated by the classes $w_{i}^{+},w_{i}^{-}$ with $i\in E$ and subject to the following relations:
  • $w_{i}^{-}=1-w_{i}^{+}$;
  • $w_{i}^{+}w_{i}^{-}=0$;
  • for any oriented circuit $C=C^{+}\sqcup C^{-}$, $\prod_{i\in C^{+}}w_{i}^{+}\prod_{j\in C^{-}}w_{j}^{-}=0$.
We will write, for $I\subseteq E$, $\omega_{I}^{+}=\prod_{i\in I}\omega_{i}^{+}$ and $\omega_{I}^{-}\prod_{i\in I}\omega_{i}^{-}$. Similarly, $\omega_{I}=\omega_{i_{1}}\omega_{i_{2}}\dots\omega_{i_{k}}$ and $\psi_{I}=\psi_{i_{1}}\psi_{i_{2}}\dots\psi_{i_{k}}$ where $I=\{i_{1},i_{2},\dots,i_{k}\}$ and the product is taken in the order induced by the ordered ground-set $E$.
Remark 3.3.
From the first and second relation it follows that $(w_{i}^{\pm})^{2}=w_{i}^{\pm}$. We consider a central circuit $C$ so that also the opposite $-C$ is a circuit. Then, using the first and the third relation for $C$, we get
$0=\prod_{i\in C^{+}}w_{i}^{+}\prod_{j\in C^{-}}(1-w_{j}^{+})=\sum_{J\subseteq C^{-}}(-1)^{|J|}w_{C^{+}}^{+}w^{+}_{J}=\sum_{J\subseteq C^{-}}(-1)^{|J|}w^{+}_{C^{+}\sqcup J}$ (2)
and for the opposite circuit
$0=\prod_{i\in C^{-}}w_{i}^{+}\prod_{j\in C^{+}}(1-w_{j}^{+})=\sum_{J\subseteq C^{+}}(-1)^{|J|}w_{J}^{+}w^{+}_{C^{-}}=\sum_{J\subseteq C^{+}}(-1)^{|J|}w^{+}_{J\sqcup C^{-}}.$ (3)
By summing the equations above with an opportune sign, we get the following relations:
$\sum_{J\subsetneq C^{-}}(-1)^{|J|}w^{+}_{C^{+}\sqcup J}-(-1)^{|C|}\sum_{J\subsetneq C^{+}}(-1)^{|J|}w^{+}_{J\sqcup C^{-}}=0$
$\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}w^{+}_{C\setminus K}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}w^{+}_{C\setminus K}=0$ (8)
From this identities, we will obtain a relation in $H^{*}(M^{a,b}(\mathcal{A}))$ by applying the map $i_{*}$.
Definition 3.4.
Let $A,B$ be disjount subsets of $E$. We denote by $\ell(A,B)$ the sign of the permutation taking the $A\sqcup B$ as a ordered subset of $E$ of $\mathcal{A}$ to the concatenation of $A$ and $B$.
Definition 3.5.
Let $A$ and $B$ be disjoint subsets of $E$ and define
$\eta_{A,B}:=(-1)^{d\ell(A,B)}\omega_{A}\psi_{B}\in H^{*}(M^{a,b}(\mathcal{A})).$
Let $B\subseteq E$ a basis of the matroid, we denote by $\operatorname{sgn}(B)$ the sign of the determinant of the matrix whose columns represent the elements $\chi_{b}$ (in a positive basis of $\Lambda$) with the order induced by $E$.
Remark 3.6.
Let $\mathcal{B}\subset\mathcal{A}$ be a sub-arrangement and $j\colon M^{a,b}_{\mathcal{A}}\hookrightarrow M^{a,b}_{\mathcal{B}}$ be the inclusion of the complements. From the definitions follow that $j^{*}\eta^{\mathcal{B}}_{C,D}=\eta_{C,D}^{\mathcal{A}}$. So in the following we will not specify the ambient space.
Lemma 3.7.
Let $\mathcal{A}$ be a totally unimodular central arrangement of rank $n$. Consider the closed immersion $i\colon M^{\mathbb{R}}(\mathcal{A})\rightarrow M^{a,b}(\mathcal{A})$ and its pushforward in cohomology $i_{*}\colon H^{0}(M^{\mathbb{R}}(\mathcal{A}))\rightarrow H^{n(a+b-1)}(M^{a,b}(\mathcal{A}))$. Let $I\subseteq E$ be an independent set and $B\subseteq E$ a basis containing $I$:
  • If $b=1$, then $i_{*}(w_{I}^{+})=\operatorname{sgn}(B)^{d}\eta_{I,B\setminus I}$.
  • If $b>1$, then $i_{*}(w_{I}^{+})=\begin{cases}0&\text{ if }\operatorname{rk}(I)\neq n,\\ \operatorname{sgn}(I)^{d}\omega_{I}&\text{ otherwise.}\end{cases}$
Proof.
Let $\mathcal{B}\subset\mathcal{A}$ be the sub-arrangement $\{H_{b}\mid b\in B\}$. Consider the inclusions $j^{\mathbb{R}}\colon M^{\mathbb{R}}(\mathcal{A})\hookrightarrow M^{\mathbb{R}}(\mathcal{B})$ and $j^{a,b}\colon M^{a,b}(\mathcal{A})\hookrightarrow M^{a,b}(\mathcal{B})$. Recall that $d=a+b-1$ is the codimension of $\mathbb{R}$ in $G$. The following diagram commutes by the naturality of pushforward in cohomology (property (3)):
[…]
The horizontal maps in the left-hand square are induced by the inclusion of the complement of $\mathcal{A}$ in the complement of $\mathcal{B}$, in $\mathbb{R}^{n}$ and $G^{n}$ respectively. Since $M^{\mathbb{R}}(\mathcal{B})\cong(\mathbb{R}\setminus 0)^{n}$ and $M^{a,b}(\mathcal{B})\cong(G\setminus e)^{n}$, the horizontal maps in the right-hand square are given by the Künneth isomorphisms. Note that the Künneth isomorphism $k^{a,b}$ depends on the sign of the basis $B$.
For the sake of notation we assume that $I$ are the first elements of $B$, the general case differs by the sign $(-1)^{d\ell(I,B\setminus I)}$ given by the reordering of $\omega_{i}$ and $\psi_{b}$ for $i\in I$ and $b\in B\setminus I$. We have
$i_{\mathcal{A}*}(\omega_{I}^{+})=i_{\mathcal{A}*}j^{\mathbb{R}*}k^{\mathbb{R}}((w^{+})^{\otimes|I|}\otimes 1^{\otimes n-|I|})=j^{a,b*}k^{a,b}((i_{*}w^{+})^{\otimes|I|}\otimes i_{*}1^{\otimes n-|I|})$
We use eq. (1) distinguishing the cases $b=1$ and $b>1$. If $b=1$,
$i_{\mathcal{A}*}(\omega_{I}^{+})=j^{a,b*}k^{a,b}(\omega^{\otimes|I|}\otimes\psi^{\otimes n-|I|})=j^{a,b*}(\operatorname{sgn}(B)^{d}\omega_{I}\psi_{B\setminus I})=\operatorname{sgn}(B)^{d}\eta_{I,B\setminus I},$
where in the last equality we use Remark 3.6. If $b>1$, $i_{*}(1)=0$ and hence ${i_{\mathcal{A}}}_{*}(w^{+}_{I})=0$ if $\operatorname{rk}(I)\neq n$. Otherwise, $I$ is a basis and
$i_{\mathcal{A}*}(\omega_{I}^{+})=j^{a,b*}k^{a,b}(\omega^{\otimes n})=j^{a,b*}(\operatorname{sgn}(I)^{d}\omega_{I})=\operatorname{sgn}(I)^{d}\omega_{I}.$
This completes the proof. ∎
Remark 3.8.
Note that, also in the case $b=1$, the image of $i_{*}$ does not depend on the choice of the (unimodular) basis $B$. In fact, if $B^{\prime}\subseteq E$ is another (unimodular) basis containing $I$, $\operatorname{sgn}(B)^{d}\psi_{B\setminus I}\omega_{I}=\operatorname{sgn}(B^{\prime})^{d}\psi_{B^{\prime}\setminus I}\omega_{I}$ by using the linear relations among the classes $\psi$ coming from the linear dependencies among the characters of $B$ and $B^{\prime}$.
Definition 3.9.
A subset $A\subseteq\mathcal{A}$ is a generalized circuit if $\operatorname{rk}(A)=|A|-1$. We denote by $C(A)$ the unique circuit contained in $A$.
Definition 3.10.
Let $C\subseteq E$ be a circuit and let $B\supset C$ such that $B\setminus{\{i\}}$ is a basis for any $i\in C$. We denote by $c_{i}$ the sign of the determinant of the matrix whose columns are the elements in $B\setminus{\{i\}}$.
Remark 3.11.
Up to a global sign, the $c_{i}$’s do not depend on the choice of $B$. The collection of these signs defines a chirotope (or equivalently an oriented matroid), cf [${}^{+}$99].
Lemma 3.12.
Let $\mathcal{A}$ be a central arrangement and $C\subseteq E$ an unimodular oriented circuit with $C=C^{+}\sqcup C^{-}$. The following relations holds in $H^{*}(M^{a,b}(\mathcal{A}))$:
  • If $b=1$,
    $\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}=0,$ (9)
    where $i_{K}\in K$.
  • If $b>1$ and $d$ odd,
    $\sum_{i\in C}(-1)^{\lvert C_{<i}\rvert}\omega_{C\setminus\{i\}}=0.$ (10)
  • If $b>1$ and $d$ even,
    $\sum_{i\in C^{-}}\omega_{C\setminus\{i\}}-\sum_{i\in C^{+}}\omega_{C\setminus\{i\}}=0.$ (11)
Proof.
Let $r$ be the rank of the arrangement $\mathcal{A}$ and $n$ be the rank of the circuit $C$. Consider the additive map given by the following composition:
[…]
where $p$ is the restriction of the projection $G^{r}\to G^{r}/\cap_{c\in C}H_{c}\simeq G^{n}$.
We consider the relation (8) in $H^{0}(M^{\mathbb{R}}(C))$ and we apply Lemma 3.7 to the essentialization of the arrangement $C$. For any independent $J\subsetneq C^{+}$ (resp. $J\subsetneq C^{-}$), the choice of $i_{K}\in K=C^{+}\setminus J$ (resp. $i_{K}\in K=C^{-}\setminus J$) corresponds to complete $C^{+}\bigsqcup J=C\setminus K$ to a basis $C\setminus i_{K}$ of the sublattice generated by the elements of $C$. By Lemma 3.7 and Remark 3.6 we have:
$p^{*}(i_{*}(\omega^{+}_{C\setminus K}))=p^{*}(c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}})=c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}},$
and so Equation (8) is mapped to
$\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta_{C\setminus K,K\setminus i_{K}}=0$
for $b=1$, and to
$\sum_{i\in C^{-}}c_{i}^{d}\omega_{C\setminus i}-\sum_{i\in C^{+}}c_{i}^{d}\omega_{C\setminus i}=0$
for $b>1$. Since $C$ is unimodular, the dependence relation among the characters of the circuit is $\sum_{i\in C}(-1)^{\lvert C_{<i}\rvert}c_{i}\chi_{i}=0$, hence $i\in C^{+}$ if and only if $c_{i}=(-1)^{\lvert C_{<i}\rvert}$. ∎
Remark 3.13.
Lemma 3.12 holds also for a central circuit $C$ in a non central arrangement $\mathcal{A}$. It is sufficient to consider the map $H^{*}(M^{a,b}(C))\rightarrow H^{*}(M^{a,b}(\mathcal{A}))$ induced by the inclusion $M^{a,b}(\mathcal{A})\hookrightarrow M^{a,b}(C)$.
Example 3.14.
{tikzpicture}{tikzpicture}
Figure 2. A picture of the arrangement in Example 3.14 for $(a,b)=(1,0)$ on the left, and $(a,b)=(0,1)$ on the right.
Let us consider the central arrangement $\mathcal{A}=\{H_{i}\}_{i=1,2,3}$ in $G^{2}$ defined by the columns of the matrix
$\begin{pmatrix}1&0&1\\ 0&1&1\end{pmatrix}$
(see Figure 2). This is an arrangement of type $A_{3}$, hence $M(\mathcal{A})$ is the configuration space of $3$ points in $G$ (c.f. Section 6). The subspace $H^{*}(G^{2})\subset H^{*}(M(\mathcal{A}))$ is generated by the classes $\psi_{1}^{j},\psi_{2}^{j}$ and $\psi_{3}^{j}=\psi_{1}^{j}+\psi_{2}^{j}$, with $j=1,\cdots,a$. Furthermore, we have defined the classes $\omega_{i}\in H^{a+b-1}(M(\mathcal{A}))$ for $i=1,2,3$. The circuit $C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\}$ gives the following relations
  • if $b=1$,
    $\omega_{1}\omega_{2}-(-1)^{d}\omega_{2}\omega_{3}-\omega_{1}\omega_{3}-\omega_{3}\psi_{1}=0,$ (12)
  • if $b>1$,
    $\omega_{1}\omega_{2}-(-1)^{d}\omega_{2}\omega_{3}-\omega_{1}\omega_{3}=0.$ (13)
4. Non-central arrangements
Now, we deal with the case of non-central totally unimodular arrangements. Firstly, we prove a variation of the Brieskorn lemma, a deletion-restriction short exact sequence, and an identity between Poincaré and characteristic polynomial. We will use these results in Section 5 for unimodular covers.
Given $H\in\mathcal{A}$ a subvariety of the arrangement, we denote by $\mathcal{A}^{\prime}:=\mathcal{A}\setminus\{H\}$ the deletion of $H$ and $\mathcal{A}^{\prime\prime}$ the restriction to $H$, i.e. the arrangement of connected components of $K\cap H$ for $K\in\mathcal{A}^{\prime}$. For any layer $p\in\mathcal{L}(\mathcal{A})$ we consider the local arrangement $\mathcal{A}_{p}:=\{H\in\mathcal{A}\mid p\subseteq H\}$ at $p$ and let $g_{p}\colon M(\mathcal{A})\to M(\mathcal{A}_{p})$ be the inclusion of the complements.
Theorem 4.1.
Let $\mathcal{A}$ be an arrangement of rank $r$ and $H\in\mathcal{A}$ a subvariety, then:
  1. (1)
    $\iota_{*}\colon H^{*}(M(\mathcal{A}^{\prime\prime}))\to H^{*+a+b}(M(\mathcal{A}^{\prime}))$ is the zero map,
  2. (2)
    the map $\oplus_{p\in\mathcal{L}^{r}}g_{p}^{*}\colon\bigoplus_{p\in\mathcal{L}^{r}}H^{*}(M(\mathcal{A}_{p}))\to H^{*}(M(\mathcal{A}))$ is surjective.
Proof.
In the central case the theorem follows from [LTY21, Theorem 7.6]. We prove the claims by induction on the cardinality of the arrangement $\mathcal{A}$. The base case is a central arrangement and the results hold. The assumption that $\mathcal{A}$ is not central implies $\lvert\mathcal{A}_{p}\rvert<\lvert\mathcal{A}\rvert$ for any $p\in\mathcal{L}^{r}$. Consider $p\in\mathcal{L}(\mathcal{A}^{\prime\prime})$ and let $\tilde{p}\in\mathcal{L}(\mathcal{A}^{\prime})$ be the unique minimal layer containing $p$ (if $H$ is not a coloop then $p=\tilde{p}$), the diagram
{tikzcd}
is a pullback diagram and so $\iota_{*}\circ g^{\prime\prime*}_{p}=g^{\prime*}_{\tilde{p}}\circ\iota_{p*}$. Since the bundary map of the long exact sequence and the Thom isomorphism are functorial, the following diagram commutes:
{tikzcd}
For $p\in\mathcal{L}(\mathcal{A})\setminus\mathcal{L}(\mathcal{A}^{\prime\prime})$ we have a similar diagram
{tikzcd}
because $\mathcal{A}^{\prime}_{p}=\mathcal{A}_{p}$. We consider the direct sum of previous sequences for all $p\in\mathcal{L}(\mathcal{A})$. Since $\lvert\mathcal{A}^{\prime\prime}\rvert<\lvert\mathcal{A}\rvert$, the map $\oplus_{p\in\mathcal{L}^{r}}g_{p}^{\prime\prime*}$ is surjective and $\iota_{p*}=0$ by inductive step:
{tikzcd}
It follows that $\iota_{*}=0$. The two rows are exact because they are long exact sequences of pairs and by diagram chasing the map $\oplus_{p\in\mathcal{L}^{r}}g_{p}^{*}$ is surjective. ∎
The corollary below follows directly from Theorem 4.1.
Corollary 4.2.
Let $\mathcal{A}$ be an arrangement of rank $r$ and $H\in\mathcal{A}$ a subvariety. The following sequence is exact
$0\to H^{*}(M(\mathcal{A}^{\prime}))\xrightarrow{j^{*}}H^{*}(M(\mathcal{A}))\xrightarrow{\operatorname{res}}H^{*-(a+b-1)}(M(\mathcal{A}^{\prime\prime}))\to 0$
Corollary 4.3.
For any abelian arrangement $\mathcal{A}$, the Poincaré polynomial only depends on the characteristic polynomial of the arrangement:
$P_{\mathcal{A}}(t)=(-t^{a+b-1})^{r}\chi_{\mathcal{A}}\Big{(}-\frac{(1+t)^{a}}{t^{a+b-1}}\Big{)}.$
Proof.
The identity follows by induction using the exact sequence in Theorem 4.1. ∎
The same formula for the Poincaré polynomial is proven for central abelian arrangements in [LTY21, Theorem 7.7]
Example 4.4.
{tikzpicture}{tikzpicture}
Figure 3. A picture of the non central arrangement in Example 4.4 when $(a,b)=(1,0)$ on the left, and $(a,b)=(0,1)$ on the right.
Let us consider the arrangement obtained from the arrangement in the example 3.14 by adding the two hyperplanes $H_{2^{\prime}}=\{\chi_{2}^{-1}(-e)\}$ and $H_{3^{\prime}}=\{(\chi_{1}+\chi_{2})^{-1}(-e)\}$ as shown in Figure 3.
The arrangement is not central anymore, but it is still unimodular. As in the previous example, the cohomology $H^{*}(M(\mathcal{A}))$ is generated by the classes $\psi_{1}^{j},\psi_{2}^{j},\psi_{3}^{j}=\psi_{1}^{j}+\psi_{2}^{j}$, $j=1,\cdots,a$ and $\omega_{1},\omega_{2},\omega_{2^{\prime}},\omega_{3},\omega_{3^{\prime}}$. There are two central circuits:
$C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\},\quad C^{\prime}=(C^{\prime})^{+}\sqcup(C^{\prime})^{-}=\{1,2^{\prime}\}\sqcup\{3^{\prime}\}$
When $b=1$, the first circuit leads to the same relation eq. (12) of Example 3.14, while the second one to the new relation:
$\omega_{1}\omega_{2^{\prime}}+(-1)^{d}\omega_{2^{\prime}}\omega_{3^{\prime}}-\omega_{1}\omega_{3^{\prime}}-\omega_{3^{\prime}}\psi_{1}=0.$ (14)
The case $b>1$ is analogous. The Poincaré polynomial of the complement is given by $P_{\mathcal{A}}(t)=(1+t)^{2a}+5(1+t)^{a}+6$. The four non-central circuits provide the following relations:
$\omega_{2}\omega_{2^{\prime}}=0,$ $\omega_{3}\omega_{3^{\prime}}=0,$
$\omega_{1}\omega_{2}\omega_{3^{\prime}}=0,$ $\omega_{1}\omega_{2^{\prime}}\omega_{3}=0,$
see eq. (26) of Theorem 5.11.
5. Non-unimodular case
The poset of layers of a non-unimodular arrangement $\mathcal{A}$ is no longer a geometric lattice, due to multiple connected components of the intersections. Indeed, the combinatoric data of $\mathcal{A}$ are encoded in a generalization of a matroid, basically a matroid together with a multiplicity function, known as arithmetic matroid and first introduced by L. Moci and M. D’Adderio in [DM13]. We deal with the non-unimodular case adapting the argument in [${}^{+}$20], by constructing an unimodular covering for an abelian arrangement $\mathcal{A}$.
Definition 5.1.
Let $\Lambda$ be a lattice and $\mathcal{A}$ an arrangement in $G^{\Lambda}$, define the multiplicity
$m(A)=[\Lambda^{A}:\Lambda_{A}]$
where $\Lambda_{A}:=\langle A\rangle\subseteq\Lambda$ and $\Lambda^{A}$ is the radical of $\Lambda_{A}$ in $\Lambda$, i.e. $\Lambda^{A}=(\mathbb{Q}\otimes_{\mathbb{Z}}\Lambda_{A})\cap\Lambda$. This number $m(A)$ is the cardinality of the torsion subgroup of $\Lambda/\Lambda_{A}$ and the multiplicity function of the matroid associated to $\mathcal{A}$.
Suppose $\mathcal{A}=\{\chi_{0},\cdots,\chi_{r}\}$ consists of one generalized circuit $X=\{0,\cdots,r\}=C\sqcup F$, with $n=\operatorname{rk}(C)$. Let
$a_{i}=\begin{cases}m(X)\prod_{j\in C\setminus\{i\}}m(C\setminus\{j\})&\text{ for }i\in C\\ m(X)&\text{ for }i\in F\end{cases}$
and let $\Lambda^{\prime}$ be the lattice in $\Lambda\otimes\mathbb{Q}$ generated by the characters $\frac{\chi_{i}}{a_{i}}$.
Lemma 5.2.
The lattice Λsuperscriptnormal-$\Lambda^{\prime}$ contains $\Lambda$ and the inclusion induces a covering $\pi\colon G^{\Lambda^{\prime}}\rightarrow G^{\Lambda}$ of degree $m(C)^{a}m(X)^{a(r-1)}\prod_{i\in C}m(C\setminus i)^{a(n-1)}$.
Proof.
The proof of the first statement is analogous to that of [${}^{+}$20, Lemma 6.4]. Let $i\in C$, as in [${}^{+}$20, Lemma 6.5] we compute the index of $\Lambda$ in $\Lambda^{\prime}$
$[\Lambda^{\prime}:\Lambda]$ $=\frac{[\Lambda^{\prime}:\Lambda_{X\setminus i}]}{[\Lambda:\Lambda_{X\setminus i}]}=\frac{\prod_{j\in X\setminus i}a_{j}}{m(X\setminus i)}=\frac{m(C)m(X)^{r}\prod_{j\in C\setminus i}\prod_{l\in C\setminus j}m(C\setminus l)}{m(X)m(C\setminus i)}$
$=m(C)m(X)^{r-1}\prod_{i\in C}m(C\setminus i)^{n-1}$
It follows that the degree of the covering is $[\Lambda^{\prime}:\Lambda]^{a}$. ∎
Let $\mathcal{A}_{U}$ be the arrangement in $U:=G^{\Lambda^{\prime}}$ given by the set of connected components of the preimages of the subvarieties in $\mathcal{A}$
$\mathcal{A}_{U}=\bigcup_{H\in\mathcal{A}}\pi_{0}(\pi^{-1}(H)).$
Lemma 5.3 ([${}^{+}$20, Lemma 6.7]).
The arrangement $\mathcal{A}_{U}$ is unimodular.
Let $H_{i}\in\mathcal{A}$, $L$ a connected component of $\pi^{-1}(H_{i})$ and $q\in L$. The subvariety $L$ has equation $\frac{\hat{\chi_{i}}}{a_{i}}=\frac{\hat{\chi_{i}}}{a_{i}}(q)$, where $\hat{\chi}=\chi\circ\pi$. We define
$\omega^{U}_{i}(q)=\bigg{(}\frac{\hat{\chi_{i}}}{a_{i}}-\frac{\hat{\chi_{i}}}{a_{i}}(q)\bigg{)}\bigg{|}_{M(\mathcal{A}_{U})}^{*}(\omega)\quad\in H^{a+b-1}(M(\mathcal{A}_{U})),$ (15)
$\psi_{i}^{j,U}=\bigg{(}\frac{\hat{\chi_{i}}}{a_{i}}\bigg{)}\bigg{|}_{M(\mathcal{A}_{U})}^{*}(\psi^{j})\quad\in H^{1}(M(\mathcal{A}_{U})),$ (16)
$\psi_{i}^{U}=\bigg{(}\frac{\hat{\chi_{i}}}{a_{i}}\bigg{)}\bigg{|}_{M(\mathcal{A}_{U})}^{*}(\psi)\quad\in H^{a}(M(\mathcal{A}_{U})).$ (17)
Remark 5.4.
For $H_{i}\in\mathcal{A}$, let $L$ be a connected component of $\pi^{-1}(H_{i})$ and $p,q\in L$. Then we have $\frac{\hat{\chi_{i}}}{a_{i}}(q)=\frac{\hat{\chi_{i}}}{a_{i}}(p)$ and hence $\omega^{U}_{i}(q)=\omega^{U}_{i}(p)$.
As in Section 3, for $I\subseteq E$, we write $\omega_{I}^{U}=\omega_{i_{1}}^{U}\omega_{i_{2}}^{U}\dots\omega_{i_{k}}^{U}$ and $\psi_{I}^{U}=\psi_{i_{1}}^{U}\psi_{i_{2}}^{U}\dots\psi_{i_{k}}^{U}$ where $I=\{i_{1},i_{2},\dots,i_{k}\}$ and the product is taken in the order induced by the ordered groundset $E$.
Let $A\subseteq E$, $W$ a connected component of $\bigcap_{i\in A}H_{i}$ and $p\in W$. Since $\pi^{*}$ is injective, we define $\omega_{W,A}$ as the unique class in $H^{*}(M(\mathcal{A}))$ such that
$\pi^{*}(\omega_{W,A})=\frac{1}{|L\cap\pi^{-1}(p)|}\sum_{q\in\pi^{-1}(p)}\omega_{A}^{U}(q)=\frac{1}{|L\cap\pi^{-1}(p)|}\sum_{\tilde{L}\text{ c.c. of }\pi^{-1}(W)}\omega_{A}^{U}(q_{\tilde{L}})$ (18)
where $L$ is any connected component of $\pi^{-1}(W)$ and $q_{\tilde{L}}\in\tilde{L}\cap\pi^{-1}(p)$ for each connected component $\tilde{L}$ of $\pi^{-1}(W)$. For any $A,B\subseteq E$ disjoint, we denote
$\eta^{U}_{A,B}(q)=(-1)^{d\ell(A,B)}\omega_{A}^{U}(q)\psi_{B}^{U}\in H^{*}(M(\mathcal{A}_{U}))$
and
$\eta_{W,A,B}=(-1)^{d\ell(A,B)}\omega_{W,A}\psi_{B}\in H^{*}(M(\mathcal{A})).$
Lemma 5.5.
Let $A,B\subseteq E$ such that $A\cup B$ is a dependent set and let $W$ be a connected component of $\bigcap_{i\in A}H_{i}$. Then, for any $j=1,\cdots,a$,
$\omega_{W,A}\psi^{j}_{B}=0.$ (19)
Proof.
By definition, it is enough to prove that, for any connected component $L$ of $\pi^{-1}(W)$, $\omega^{U}_{A}(q_{L})\psi_{B}^{j,U}=0$ in $H^{*}(M(\mathcal{A}_{U}))$. This follows from $(\psi_{i}^{j,U})^{2}=0$, $\omega_{i}^{U}(q_{L})\psi_{i}^{j,U}=0$ and from the linear dependences between $\{\chi_{i},i\in A\cup B\}$. ∎
Lemma 5.6.
Let $X$ be a generalized circuit and $A,B\subseteq X$ such that $A\sqcup B$ is a maximal independent set. Let $W$ be a connected component of $\bigcap_{i\in A}H_{i}$ and $p\in W$. Then,
$\pi^{*}(\eta_{W,A,B})=\frac{m(A\cup B)^{a}}{m(A)^{a}}\sum_{q\in\pi^{-1}(p)}\eta_{A,B}^{U}(q)$
Proof.
By definition,
$\pi^{*}(\psi_{B})=\big{(}\prod_{i\in B}a_{i}^{a}\big{)}\psi_{B}^{U}$
The computation of $|L\cap\pi^{-1}(p)|$ is analogous to that in Lemma 6.8 in [${}^{+}$20]. The cardinality of the preimage of $p$ is the degree of the covering. The number of connected components of $\pi^{-1}(W)$ is
$([\Lambda(A):\Lambda^{A}])^{a}=\bigg{(}\frac{\prod_{i\in A}a_{i}}{m(A)}\bigg{)}^{a}$
Therefore,
$|L\cap\pi^{-1}(p)|$ $=\frac{m(C)^{a}m(X)^{a(r-1)}\prod_{i\in C}m(C\setminus i)^{a(n-1)}m(A)^{a}}{\prod_{i\in A}a_{i}^{a}}$
$=\frac{m(C)^{a}m(X)^{a(r-1)}\prod_{i\in C}m(C\setminus i)^{a(n-1)}m(A)^{a}}{m(X)^{a|A|}\prod_{i\in A\cap C}\prod_{j\in C\setminus i}m(C\setminus j)^{a}}$
$=m(C)^{a}m(X)^{a(r-|A|-1)}m(A)^{a}\prod_{i\in C}m(C\setminus i)^{a(n-|A\cap C|-1)}\prod_{i\in A\cap C}m(C\setminus i)^{a}$
It follows that
$\pi^{*}(\eta_{W,A,B})=\frac{\prod_{i\in B}a_{i}^{a}}{|L\cap\pi^{-1}(p)|}(-1)^{d\ell(A,B)}\sum_{q\in\pi^{-1}(p)}\omega_{A}^{U}(q)\psi_{B}^{U}$
where,
$\frac{\prod_{i\in B}a_{i}^{a}}{|L\cap\pi^{-1}(p)|}$ $=\frac{m(X)^{a|B|}\prod_{i\in C}m(C\setminus i)^{a|B\cap C|}}{\prod_{i\in B\cap C}m(C\setminus i)^{a}}\cdot\frac{1}{|L\cap\pi^{-1}(p)|}$
$=\frac{m(X)^{a(|A|+|B|-r+1)}\prod_{i\in C}m(C\setminus i)^{a(|B\cap C|+|A\cap C|-n+1)}}{m(C)^{a}m(A)^{a}\prod_{i\in A\cap C}m(C\setminus i)^{a}\prod_{i\in B\cap C}m(C\setminus i)^{a}}$
$=\frac{m(X)^{a}\prod_{i\in C}m(C\setminus i)^{a}}{m(C)^{a}m(A)^{a}\prod_{i\in(A\cup B)\cap C}m(C\setminus i)^{a}}$
$=\frac{m(X)^{a}m((A\cup B)\cap C)^{a}}{m(C)^{a}m(A)^{a}}=\frac{m(A\cup B)^{a}}{m(A)^{a}}$
In this last computation the final equality comes from property $(3)$ of arithmetic matroids, see [DM13]. ∎
Lemma 5.7.
Let $X=C\sqcup F$ be a generalized oriented circuit, $Y$ be a connected component of $\bigcap_{i\in X}H_{i}$, then
  • If $b=1$,
    $\begin{split}\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\frac{m(X\setminus K)^{a}}{m(X\setminus i_{K})^{a}}\eta_{W,X\setminus K,K\setminus i_{K}}&\\ -\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\frac{m(X\setminus K)^{a}}{m(X\setminus i_{K})^{a}}\eta_{W,X\setminus K,K\setminus i_{K}}&=0,\end{split}$ (20)
    where $i_{K}\in K$ and, for each summand, $W$ is the connected component of $\bigcap_{i\in X\setminus K}H_{i}$ such that $Y\subseteq W$.
  • If $b>1$ and $d$ odd,
    $\sum_{i\in C}(-1)^{|C_{<i}|}\omega_{Y,X\setminus i}=0.$ (21)
  • If $b>1$ and $d$ even,
    $\sum_{i\in C^{-}}\omega_{Y,X\setminus i}-\sum_{i\in C^{+}}\omega_{Y,X\setminus i}=0.$ (22)
Proof.
We apply Lemma 3.12 to the circuit $C$ in the abelian variety $U=G^{\Lambda^{\prime}}$. Let $p\in Y$ be any point. If $b=1$, we have:
$\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{C\setminus K,K\setminus i_{K}}(q)-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{C\setminus K,K\setminus i_{K}}(q)=0$
for all $q\in\pi^{-1}(p)$. We multiply this equation by $\omega^{U}_{F}(q)$, and we get:
$\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{k}}(q)=0.$
Summing over all $q\in\pi^{-1}(p)$, we obtain:
$0=$ $\sum_{q\in\pi^{-1}(p)}\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)-\sum_{q\in\pi^{-1}(p)}\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)$
$=$ $\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\sum_{q\in\pi^{-1}(p)}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\sum_{q\in\pi^{-1}(p)}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)$
$=$ $\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\frac{m(X\setminus K)^{a}}{m(X\setminus i_{K})^{a}}\pi^{*}(\eta_{W,X\setminus K,K\setminus i_{K}})$
$-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\frac{m(X\setminus K)^{a}}{m(X\setminus i_{K})^{a}}\pi^{*}(\eta_{W,X\setminus K,K\setminus i_{K}}).$
Since $\pi^{*}$ is injective, equation (20) follows.
If $b>1$, through an analogous computation, equations (21) and (22) hold.
Example 5.8.
{tikzpicture}{tikzpicture}
Figure 4. A picture of the non unimodular arrangement in Example 5.8 when $(a,b)=(1,0)$ on the left, and $(a,b)=(0,1)$ on the right.
Let us consider the arrangement $\mathcal{A}=\{H_{i}\}_{i=1,2,3,4}$ in $G^{3}$ associated to the matrix
$\begin{pmatrix}1&0&2&1\\ 0&1&1&0\\ 0&0&0&2\end{pmatrix}$
(see Figure 4). The generalized circuits are $C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\}$ and $X=C\sqcup F=\{1,2,3\}\sqcup\{4\}$.
Let $W$ be the layer $H_{1}\cap H_{2}\cap H_{3}$. If $b>1$, applying formula (20) to the central circuit $C$ we obtain the relation
$\omega_{W,12}+(-1)^{d}\omega_{W,23}-\omega_{W,13}+\frac{1}{2^{a}}\omega_{3}\psi_{2}=0.$ (23)
Notice that in the case $a=1$ the unimodular covering of $\mathcal{A}$ is the arrangement in Example 4.4, and the pullback of eq. (23) is the sum of equations (12) and (14).
Now consider the generalized circuit $X$, we get $2^{a}$ relations because the intersection $H_{1}\cap H_{2}\cap H_{3}\cap H_{4}=W\cap H_{4}$ has $2^{a}$ distinct connected components, $p_{1},\cdots,p_{2^{a}}$ and for any $i=1,\cdots,2^{a}$:
$\omega_{p_{i},124}+(-1)^{d}\omega_{p_{i},234}-\omega_{p_{i},134}-\frac{1}{4^{a}}\omega_{Z,34}\psi_{2}=0,$ (24)
where $Z$ is the layer $H_{3}\cap H_{4}$. The Poincaré polynomial of the complement is $(t+1)^{3a}+4(t+1)^{2a}+7(t+1)^{a}+6$.
Lemma 5.9.
The classes $\omega_{W,A}$ for all independent sets $A\subseteq E$ and all connected components $W$ of $\bigcap_{i\in A}H_{i}$, generate $H^{*}(M(\mathcal{A}))$ as a $H^{*}(G^{r})$-module.
Proof.
Fix $i\in E$, we proceed by induction using the exact sequence (of $\mathbb{Z}$-module) of Corollary 4.2 given by deletion and contraction with respect to $H_{i}$:
$0\to H^{*}(M(\mathcal{A}^{\prime}))\xrightarrow{j^{*}}H^{*}(M(\mathcal{A}))\xrightarrow{\operatorname{res}}H^{*-(a+b-1)}(M(\mathcal{A}^{\prime\prime}))\to 0$
Let $\mathcal{K}\subseteq H^{*}(M(\mathcal{A}))$ be the $H^{*}(G^{r})$-submodule generated by all $\omega_{W,A}$, $\mathcal{K}^{\prime}\subseteq H^{*}(M(\mathcal{A}^{\prime}))$ the $H^{*}(G^{r})$-submodule generated by all $\omega_{W,A}^{\prime}$ with $i\not\in A$, and $\mathcal{K}^{\prime\prime}\subseteq H^{*}(M(\mathcal{A}^{\prime\prime}))$ the $H^{*}(G^{r-1})$-submodule generated by all $\omega_{W,A}^{\prime\prime}$ with $i\in A$.
By definition of the classes $\omega_{W,A}$ we have $j_{*}(\omega_{W,A}^{\prime})=\omega_{W,A}$ and
$\operatorname{res}(\omega_{W,A})=\begin{cases}\omega^{\prime\prime}_{W,A\setminus\{i\}}&\textnormal{if }i\in A,\\ 0&\textnormal{otherwise.}\end{cases}$
Since the pullback map in cohomology $H^{*}(G^{r})\twoheadrightarrow H^{*}(G^{r-1})$ is surjective and the residue map is a morphism of $H^{*}(G^{r})$-modules, we have
{tikzcd}
By diagram chasing the inductive step follows, this completes the proof. ∎
Definition 5.10.
Let $R$ be the free $H^{*}(G^{r})$-module generated by the classes $\omega_{W,A}\in H^{*}(M(\mathcal{A}))$ with $A\subseteq E$ independent and $W$ connected component of $\bigcap_{i\in A}H_{i}$. We endow $R$ with a ring structure by defining the multiplication
$\omega_{W,A}\omega_{W^{\prime},A^{\prime}}=\begin{cases}0\quad\text{ if }A,A^{\prime}\text{ are not disjoint or }A\sqcup A^{\prime}\text{ dependent,}\\ (-1)^{d\ell(A,A^{\prime})}\sum_{L\text{ c.c. of}W\cap W^{\prime}}\omega_{L,A\sqcup A^{\prime}}\quad\text{ otherwise.}\end{cases}$ (25)
Notice that, for any $A,B\subseteq E$, and for any connected component $W$ of $\bigcap_{i\in A}H_{i}$, $\eta_{W,A,B}\in R$.
Theorem 5.11.
Let $\mathcal{A}$ be an arrangement in $G^{r}$, where $G=\mathbb{R}^{b}\times(S^{1})^{a}$. The integer cohomology of the complement $H^{*}(M(\mathcal{A});\mathbb{Z})$ is the quotient of $R$ by the following relations:
  • Whenever $A\sqcup B$ and $A^{\prime}\sqcup B^{\prime}$ are not disjoint or $A\sqcup B\sqcup A^{\prime}\sqcup B^{\prime}$ is dependent,
    $\eta_{W,A,B}\eta_{W^{\prime},A^{\prime},B^{\prime}}=0$ (26)
  • For any generalized circuit $X=C\sqcup F$ with $C=C^{+}\sqcup C^{-}$, and for any connected component $Y$ of $\bigcap_{i\in X}H_{i}$,
    • If $b=1$,
      $\begin{split}\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}&(-1)^{|K|}c_{i_{K}}^{d}\frac{m(X\setminus K)^{a}}{m(X\setminus i_{K})^{a}}\eta_{W,X\setminus K,K\setminus i_{K}}\\ &-\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\frac{m(X\setminus K)^{a}}{m(X\setminus i_{K})^{a}}\eta_{W,X\setminus K,K\setminus i_{K}}=0,\end{split}$ (27)
      for some $i_{K}\in K$ and where, for each summand, $W$ is the connected component of $\bigcap_{i\in X\setminus K}H_{i}$ such that $Y\subseteq W$.
    • If $b>1$ and $d$ odd,
      $\sum_{i\in C}(-1)^{|C_{<i}|}\eta_{Y,X\setminus i,\emptyset}=0.$ (28)
    • If $b>1$ and $d$ even,
      $\sum_{i\in C^{-}}\eta_{Y,X\setminus i,\emptyset}-\sum_{i\in C^{+}}\eta_{Y,X\setminus i,\emptyset}=0.$ (29)
Proof.
The cohomology ring is a quotient of $R$ by Lemma 5.9. Relation (26) holds by Lemma 5.5 and relations (27), (28) by Lemma 5.7.
Let $I$ be the ideal given by relations (26)-(28). There is a surjective map $R/I\rightarrow H^{*}(M(\mathcal{A}))$. We prove that this map is an isomorphism by induction on the cardinality of $\mathcal{A}$. The base case is trivial.
Let us fix $i\in E$. Let $R^{\prime}$, $R^{\prime\prime}$, $I^{\prime}$ and $I^{\prime\prime}$ be defined analogously with respect to the deletion $\mathcal{A}^{\prime}$ and the contraction $\mathcal{A}^{\prime\prime}$ respectively. By induction hypothesis $R^{\prime}/I^{\prime}\cong H^{*}(M(\mathcal{A}^{\prime}))$ and $R^{\prime\prime}/I^{\prime\prime}\cong H^{*}(M(\mathcal{A}^{\prime\prime}))$. Let us consider the following diagram:
{tikzcd}
where:
    the first horizontal map is defined by the inclusions $R^{\prime}\subseteq R$ and $I^{\prime}\subseteq I$;
    the second horizontal map is induced by the map $g\colon R\rightarrow R^{\prime\prime}$ given by
    $\omega_{W,A}z\mapsto\begin{cases}\omega_{W,A\setminus i}\bar{z},&\text{ if }i\in A,\\ 0&\text{ otherwise,}\end{cases}$
    for any $z\in H^{*}(G^{r})$, where $\bar{z}$ is the image of $z$ through $r\colon H^{*}(G^{r})\rightarrow H^{*}(H_{i})$. The map is well defined since the image of $I$ is $I^{\prime\prime}$;
    the second row is the deletion-contraction exact sequence of Corollary 4.2.
The left square of diagram commutes because $j^{*}(\eta^{\prime}_{W,A,B})=\eta_{W,A,B}$, as the right one, since $\operatorname{res}(\eta_{W,A,B})=\eta_{W,A\setminus\{i\},B}^{{}^{\prime\prime}}$ if $i\in A$ and $\operatorname{res}(\eta_{W,A,B})=0$ otherwise. This implies that the first horizontal map is injective and the second one surjective. We now want to prove that the central vertical map is injective.
Firstly, observe that for any relation $r$ of type (27) and any generator $\eta_{W,A,B}$ the element $r\eta_{W,A,B}$ is linear combination of relations of types (26) and (27). In particular, the ideal $I$ is generated by (26) and (27) as $\mathbb{Z}$-module. It follows that the map $g_{|I}\colon I\twoheadrightarrow I^{\prime\prime}$ is surjective.
[…]
6. Configuration spaces
In this section, we apply the main result of this article to the cohomology algebra of ordered configuration spaces in $\mathbb{R}^{b}\times(S^{1})^{a}$.
Let $X$ be a topological space, the configuration space of $n$ points in $X$ is:
$\operatorname{Conf}_{n}(X)=\{(x_{1},\dots,x_{n})\in X^{n}\mid x_{i}\neq x_{j}\,\forall i\neq j\}.$
As above, let $G=\mathbb{R}^{b}\times(S^{1})^{a}$, and consider the totally unimodular and central arrangement
$\mathcal{A}_{n}=\{H_{ij}\}_{\begin{subarray}{c}i,j\in[n]\\ i<j\end{subarray}},\quad H_{ij}=(\chi_{i}-\chi_{j})^{(-1)}(e),$
where $\chi_{i}\colon G^{n}\rightarrow G$ is the projection on the $i$-th component. Notice that
$\operatorname{Conf}_{n}(G)=M^{a,b}(\mathcal{A}_{n}).$
Cohen and Taylor [CT78] used a spectral sequence to compute the cohomology of configuration spaces in $\mathbb{R}^{b}\times M$, for $b\geq 1$ and $M$ a connected manifold (see [CT78, Example 1]). More precisely, there exists a filtration $F_{*}$ on $H^{*}(\operatorname{Conf}_{n}(\mathbb{R}^{b}\times M);\mathbb{K})$ for a field $\mathbb{K}$ and they described explicitly the associated graded
$\operatorname{gr}_{F_{*}}H^{*}(\operatorname{Conf}_{n}(\mathbb{R}^{b}\times M);\mathbb{K})$
as a ring. In the case $M=(S^{1})^{a}$ and $b>1$, Theorem 5.11 implies that
$\operatorname{gr}_{F_{*}}H^{*}(\operatorname{Conf}_{n}(G);\mathbb{K})\simeq H^{*}(\operatorname{Conf}_{n}(G);\mathbb{K})$
as ring. In the case $b=1$, we have the opposite behaviour: indeed, the two rings $\operatorname{gr}_{F_{*}}H^{*}(\operatorname{Conf}_{n}(G;\mathbb{K}))$ and $H^{*}(\operatorname{Conf}_{n}(G);\mathbb{K})$ are not canonically isomorphic. Finally, our result extend the coefficients from fields to integers.
Now, we write down the presentation of $H^{*}(\operatorname{Conf}_{n}(G);\mathbb{Z})$ by using Theorem 5.11. The ground set of the associated matroid is the set $E=\{ij|i,j\in[n],i<j\}$ with the lexicographic order. The poset of layers is $\mathcal{L}(\mathcal{A}_{n})=\Pi_{[n]}$ the poset of partitions of $[n]$. The circuits of the arrangement are of the form $\{ij,ik,jk\}$ with $i,j,k\in[n]$ and $i<j<k$.
By Theorem 5.11, the cohomology of the configuration space is generated as a $H^{*}(G^{n})$-module by the classes $\omega_{ij}\in H^{d}(\operatorname{Conf}_{n}(G))$ with $ij\in E$.
Relations (26)-(28) become
  • $\omega_{ij}\psi_{ij}=0$, $\forall ij\in E$;
  • for $b=1$,
    $\omega_{ij}\omega_{jk}-\omega_{ij}\omega_{ik}+\omega_{jk}\omega_{ik}-\psi_{ij}\omega_{ik}=0;$ (30)
  • for $b>1$,
    $\omega_{ij}\omega_{jk}-\omega_{ij}\omega_{ik}+\omega_{jk}\omega_{ik}=0.$ (31)
Recall that in $H^{*}(G^{n})$, $\psi_{ik}=\psi_{ij}+\psi_{jk}$ and notice that relation (26) $\omega_{ij}\omega_{jk}\omega_{ik}=0$ follows by multiplying equation (30) (resp. eq. (31)) by $\omega_{ik}$. In the case $b=1$, correction terms appear in our formulas. In the article [CT78] and in this one, the element $\omega_{ij}$ represent the cohomological class of $H_{ij}$ and hence there is no canonical morphism between the two algebras $\operatorname{gr}_{F_{*}}H^{*}(\operatorname{Conf}_{n}(G;\mathbb{K}))$ and $H^{*}(\operatorname{Conf}_{n}(G);\mathbb{K})$.
Acknowledgements
R. Pagaria and M. Pismataro are partially supported by PRIN ALgebraic and TOPological combinatorics (ALTOP) n° 2022A7L229. The authors are member of the INDAM group GNSAGA, that has partially supported R. Pagaria and M. Pismataro.